OpenStudy (anonymous):

sec squared /sec squared -1=csc squared

OpenStudy (freckles):

\[\frac{\sec^2(x)}{\sec^2(x)-1}=\csc^2(x) ?\]

OpenStudy (jdoe0001):

\(\bf \cfrac{sec^2(\theta)}{sec^2(\theta)}-1=csc^2(\theta) ?\)

OpenStudy (anonymous):

yes freckles